OpenStudy (rainbow_rocks03):

@Nnesha will you help me . The simple interest on an investment of $4000 for one year is $140. What is the interest rate? A. 0.035% B. 3.5% C. 35% D. 350%

Nnesha (nnesha):

divide number of year /investment and then multiply by 100

OpenStudy (jdoe0001):

\(\large { \qquad \textit{Simple Interest Earned}\\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\to &\$140\\ P=\textit{original amount deposited}\to& \$4000\\ r=rate\to r\%\to \frac{r}{100}\\ t=years\to &1 \end{cases} }\) solve for "r"

OpenStudy (rainbow_rocks03):

so 4000/140*100 ?

OpenStudy (rainbow_rocks03):

is that correct?

OpenStudy (nothingwasthesame):

yea

OpenStudy (rainbow_rocks03):

so a or d because when i did that I got a crazy number

Nnesha (nnesha):

what's that ?

OpenStudy (rainbow_rocks03):

what's what?

Nnesha (nnesha):

crazy number ?

OpenStudy (rainbow_rocks03):

this is the number i got when i did it look. 4000/140*100=2857.142857142857

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @jdoe0001 \(\large { \qquad \textit{Simple Interest Earned}\\ I = Prt\qquad \begin{cases} I=\textit{interest earned}\to &\$140\\ P=\textit{original amount deposited}\to& \$4000\\ r=rate\to r\%\to \frac{r}{100}\\ t=years\to &1 \end{cases} }\) solve for "r" \(\color{blue}{\text{End of Quote}}\) use this formula A = 140 so replace i by 140 p by 4000

Nnesha (nnesha):

one year so t=1 substitute values into the formula

Nnesha (nnesha):

 substitute